Knowledge Check 01

Expected Activity

Activities and Activity Measures

Estimated Overhead Costs Product 1 Product 2
Hand-processing (direct labor-hours) $ 320,000 45,000 55,000
Setting up equipment (number of setups) 150,000 9,000 21,000
Administering parts inventories (number of part types) 30,000 110 90
$ 500,000

Transformation Corporation has provided the information set forth above regarding the overhead to be allocated between the two products it produces: CDE and EFG. What is the activity rate for the hand-processing units activity?
Answer #1

Activity rate for hand processing = estimated overhead / total labour hours

Total labour hours = 45000+55000= 100000

Thus, activity rate for hand processing = 320000/100000

Activity rate = $3.20 per labour hour.